# Practitioner Motives to Use Different Hyperparameter Optimization Methods

### v3

@@ 3.1. Semi-structured Interviews with Machine Learning Experts
To identify practitioners’ goals pursued in HPO and to understand decisions for specific HPO methods to achieve these goals, we chose an exploratory, qualitative research approach and conducted semi-structured expert interviews.
Data Gathering
@@
The ML experts were associated with thirteen different organizations and had an average work experience in ML of about five years (see Table 1). Ten participants had a skill level of ML innovators—practitioners engaged in ML research, including studies on core algorithms and the application of ML in scientific discoveries (Xin et al., 2021)—while the other ten were ML engineers—experienced ML practitioners with formal training in applied ML (Xin et al., 2021). The study participants used HPO in the context of computer vision, natural language processing, reinforcement learning, and time series forecasting. Interviewees from the industry also mentioned to have used these ML techniques in bioinformatics (2), robotics (2), e-commerce (1), and finance (1). Because AutoML, including HPO based on programmatic methods, aims to be domain agnostic (Feurer and Hutter, 2019), usage of AutoML tools should be independent of the different domains.
@@
We developed an interview guide for the semi-structured interviews (Myers and Newman, 2007; Gorden, 1975). The interview guide structured the interviews into four sections: briefing, HPO in ML, participant background and personal experiences, and debriefing. Each section included details on its purpose and questions to be asked in the interview. We sent the interview guide to participants in advance to help them prepare for the interview. Since the interviews focused on a past ML project, we asked the participants to choose one to discuss in their interviews. As we also recruited a few participants from personal contacts, we aimed to avoid social desirability bias; we ensured that no interviewee was acquainted with the interviewer. Moreover, we phrased the questions as neutral and open-ended to encourage unbiased and detailed responses. For each interview, we created a non-judgemental atmosphere and clarified that there are only valid answers to our questions with valuable insights important to our research. We also explained to the interviewees that their responses would be anonymized and no links to their identity would be possible.
@@
Three of the authors coded each interview together. After familiarizing themselves with the transcripts (Step 1), the authors independently coded them (Step 2) to identify HPO methods applied by practitioners, to extract practitioners’ goals in HPO, and to reveal contextual factors that influence the interviewees’ decisions for HPO methods. We incorporated contextual factors to better understand influences on practitioners’ decisions for using HPO methods. The three authors independently read the transcripts, identified quotes relevant to this study, and labeled the quote with a name (i.e., a code) that expresses a potentially relevant HPO method, goal, or contextual factor. Each of the three authors coded the transcripts independently without a predefined coding scheme. After coding each interview independently, the authors presented their coding results to each other and discussed and harmonized their results in multiple iterations per interview and across interviews. Per iteration, the analysts aimed for mutual exclusiveness of the codes while maintaining exhaustiveness. Through this explorative approach involving three analysts per interview, we aimed to reduce biases arising opinions of single analysts. The first coding iteration revealed 241 preliminary codes.
@@
After coding transcripts of 13 of the 20 interviews, the analysis of the remaining 7 transcripts did not reveal additional HPO methods, goals, and contextual factors. We assume to have reached theoretical saturation (Fusch and Ness, 2015; Guest et al., 2006) after the first 13 interviews with 7 interviews confirming this assumption. Given the heterogeneity of the interviewees and the large fraction of confirmation interviews, we assumed the results from the interviews were sufficiently robust and moved on with an online survey to collect evidence for the external validity of the coding results.
@@ 3.2. Online Survey
The online questionnaire was structured into three sections: Introduction, HPO Methods and Goals, and Contextual Factor Integration. In the Introduction section, we described the motivation for and the structure of the questionnaire. In Methods and Goals, we showed participants a matrix that listed all goals and HPO methods identified in the interview study. Participants were asked to select all pairs of HPO methods and goals to achieve specific goals. Because the main purpose of the survey was to collect evidence for the validity of the results from the interview study, study participants could only select HPO methods (i.e., manual tuning, grid search, random search, Bayesian optimization, and evolutionary algorithms) and goals (e.g., increase ML model understanding, decrease necessary computations, and decrease practitioner effort) extracted from the previous interview study. In a second question, we asked participants to indicate, for each selected pair of HPO method and goal, whether they feel to have successfully achieved each goal.
@@
To solicit participants for the online questionnaire, we contacted practitioners via email and promoted the study via social media platforms. We did not invite participants from the interview study. A total of 49 participants completed the questionnaire. Most participants worked in large organizations with more than 500 employees, including automotive companies, companies specializing in IT support and services, and universities. Table 2 shows more demographic details about the participants who completed the questionnaire.
@@
By analyzing the responses to the online questionnaire, we sought to learn how frequently practitioners tend to choose which HPO methods to pursue specific goals, given which contextual factors. We extracted the number of identical responses and related them to each other.
@@ 4. Results
The study participants (i.e., the interviewees and the survey participants) applied five principal HPO methods to achieve six goals under consideration of fourteen contextual factors (see sections 4.1.14.1.3). The study participants pursued their goals with different self-reported success (see section  4.2).
@@ 4.1.1. Hyperparameter Optimization Methods Used by Practitioners
The study participants applied five HPO methods: manual tuning, grid search, random search, Bayesian optimization, and evolutionary algorithms. Each HPO method was discussed by at least two and at most eight interviewees. Most interviewees discussed one or two HPO methods in detail. [^2] We asked the interviewees to discuss a prominent recent example of how they used HPO methods, not to discuss all HPO methods they ever used. Most survey participants primarily used manual tuning, followed by grid search, Bayesian optimization, random search, and evolutionary algorithms (see Figure 1a). Participants from academy tended to use more manual tuning and grid search than participants from industry. Most survey participants stated to have used at least three HPO methods in their past ML projects; about 1.5% have even used at least five HPO methods. Only roughly 15% of the survey participants have used only a single HPO method (see Figure 1b).
@@
Even though literature indicates that Bayesian optimization yields better results in a shorter time than evolutionary algorithms, grid search, and random search (e.g., (Snoek et al., 2012; Turner et al., 2021)), practitioners tend to use seemingly inferior HPO methods. Practitioners appear to not only aim at finding hyperparameter configurations for optimal ML model performance but also pursue different goals.
@@ 4.1.2. Goals of Practitioners Pursued with Different HPO Methods
The goal comply with target audience refers to the alignment of individuals’ behaviors with expectations of a target audience. Three interviewees from academia stated they had decided on HPO methods to comply with the expectations of their target audiences regarding applied HPO methods and the resulting ML model. For example, two of the three interviewees described Bayesian optimization as uncommon in their research communities and felt obliged to explain it in their scientific publications. This would have required additional explanations of Bayesian optimization, even though the authors assumed the exact HPO method was not relevant to their scientific work. Therefore, they decided to use grid search as they assumed this HPO method to be well-known in their research communities.
@@
Extensive searches for optimal hyperparameter configurations in large search spaces usually require a vast amount of compute resources. Necessary computations for HPO can be decreased by using an HPO method that requires fewer compute resources than other methods but is still sufficiently useful.
@@
In contrast to the previously mentioned goal, decreasing necessary computations is not self-contained. Instead, it is a supplemental goal pursued in combination with at least one of the previously introduced goals, and to cope with contextual factors. In the latter case, for example, practitioners opt to decrease necessary computations to use scarce available compute resources to sufficiently increase ML model performance.
@@
Practitioners choose HPO methods to reduce overhead, for example, in terms of the additional time required to understand the HPO method or to integrate the HPO method into workflows. One interviewee from industry stated that the overwhelming amount of different advanced HPO tools available made it difficult to choose the one most appropriate for their use case. Seemingly being subject to the paradox of choice (Schwartz, 2004), they did not feel comfortable with selecting an HPO tool. Instead, they used manual tuning. To decrease their efforts in HPO, the interviewees applied grid search and manual tuning. In particular, practitioners stated to have applied manual tuning to avoid efforts related to setting up HPO tools in cluster infrastructures.
@@
Like reducing computational requirements, decreasing practitioner effort is not self-contained, and commonly used in combination with other goals. For example, study participants used HPO tools to enhance ML model understanding by iteratively refining manually defined sets of hyperparameter values. Those practitioners leveraged HPO tools to automate the reconfiguration of defined hyperparameter sets to accelerate manual tuning.
@@
Increasing ML model understanding refers to reaching the state where a practitioner can predict changes in an ML model’s behavior due to tuning hyperparameter values based on an understanding of the ML model’s inner workings. To increase their understanding of ML models, the interviewees reported having applied manual tuning. The interviewees claimed that manual tuning can improve their understanding of hyperparameter influences on ML models as they formulate hypotheses about hyperparameter influences on ML models and evaluate them immediately. The interviewees explained that they iteratively improve their ML model understanding by tuning hyperparameter values and testing their hypotheses. This dynamic between ML practitioner and ML model is researched as interactive ML. Interactive ML supports practitioners in actively exploring the model space and testing hypotheses about hyperparameter effects (Jiang et al., 2019). In addition, visual analytics tools help practitioners by providing graphical representations of model behavior and parameter interactions, e.g.,  (Spinner et al., 2019; Heyen et al., 2020).
@@
The survey results indicate that all goals extracted from the interview study are also pursued by the survey participants (see Figure 2). More than 96% of the survey participants pursued the goal increase ML model performance and 78% aimed to achieve decrease necessary computations. 77% sought to decrease practitioner effort. 69% of the practitioners aimed to increase ML model understanding. The least pursued goals are satisfy requirements (65%) and comply with target audience (55%).
@@
Figure 3 shows how often the survey participants used HPO methods to reach the goals identified in the interview study. 69% of the survey participants tried to decrease necessary computations by using Bayesian optimization and 58% using evolutionary algorithms. About 50% of the participants tried to decrease the necessary computations by applying manual tuning. Random search and grid search were least often used, with 44% and 32%, respectively. Decreasing practitioner effort was of interest for less than 50% of the participants, with very similar responses for all HPO methods (yet a notable exception of manual tuning with only 30%). The survey participants mainly used manual tuning to better comprehend ML models. Interestingly, participants also tried to use Bayesian optimization twice as often as grid search and random search to achieve this goal despite its black-box nature. Participants from academia have only tried to satisfy requirements or be compliant with target audience half as often as participants from industry.
@@
Practitioners mostly aimed to satisfy requirements using evolutionary algorithms (42%), Bayesian optimization (36%), and manual tuning (36%). The study participants less often used grid search (25%) and random search (21%) to achieve this goal.
Comply with target audience was the least pursued goal for the participants. About 40% of the participants used evolutionary algorithms for this goal. We could not uncover differences between manual tuning and Bayesian optimization, with 30% of the participants using the respective method. Only random search and manual tuning were very seldom used to achieve this goal (about 13%).
In contrast to academia, participants from industry used seldomly Bayesian optimization or evolutionary algorithms for increasing model understanding. Participants from academia prefer to use grid search to reduce their effort. Finally, participants from industry tend to use manual tuning more for satisfying requirements.
Apparently, participants use different HPO methods even though they aim to achieve similar goals. However, a clear mapping of HPO methods to goals is hardly possible. For example, one interviewee chose manual tuning to increase ML model understanding while another interviewee preferred Bayesian optimization.
@@
Due to the ambiguities of responses, the selection of HPO methods appears not to be simply explainable based on practitioners’ goals. For example, practitioners used different HPO methods to reach identical goals. This sheds a dim light on why practitioners used what HPO method. To better understand practitioners’ reasons for selecting HPO methods, contextual factors that influence practitioners in their decisions for HPO methods to reach goals seem important. By also illuminating contextual factors in combination with the use of HPO methods and goals to be achieved, the motives of practitioners to use HPO methods should be better understood.

### v4

@@ 3.1. Semi-structured Interviews with Machine Learning Experts
To identify practitioners’ goals pursued in HPO and understand decisions for specific HPO methods to achieve these goals, we chose an exploratory, qualitative research approach and conducted semi-structured expert interviews.
Data Collection
@@
The interviewed experts were all ML practitioners—individuals who regularly develop ML models for practical use in research or industry. All interviewees had actively contributed to at least one successfully deployed ML model in research or industry. The experts were associated with thirteen different organizations and had an average work experience in ML of about five years (see Table 1). Among the interviewees, two held a PhD and 12 were PhD students from academia with Master’s degrees—most nearing completion of their doctorates. All industry participants held at least a Master’s degree and had, on average, more than five years of experience developing ML models for production. The 12 PhD students worked at universities in the field of applied ML. One PhD worked in the industry as an ML engineer, and another worked at a university in the field of applied ML. Two interviewees with Bachelor’s degrees were pursuing Master’s degrees while working as ML developers in industry as student trainees.
Ten participants had a skill level of ML innovators—practitioners engaged in ML research, including studies on core algorithms and the application of ML in scientific discoveries (Xin et al., 2021)—while the other ten were ML engineers—experienced ML practitioners with formal training in applied ML (Xin et al., 2021). The study participants used HPO in the context of computer vision, natural language processing, reinforcement learning, and time series forecasting. Interviewees from the industry also mentioned to have used these ML techniques in bioinformatics (2), robotics (2), e-commerce (1), and finance (1). Because AutoML, including HPO based on programmatic methods, aims to be domain agnostic (Feurer and Hutter, 2019), usage of AutoML tools should be independent of the different domains.
@@
We developed an interview guide for the semi-structured interviews (Myers and Newman, 2007; Gorden, 1975). The interview guide structured the interviews into four sections: briefing, HPO in ML, participant background and personal experiences, and debriefing. Each section outlined its purpose and the corresponding interview questions. We sent the interview guide to participants in advance to help them prepare for the interview. Since the interviews focused on a past ML project, we asked the participants to choose one to discuss in their interviews. As we also recruited a few participants from personal contacts, we aimed to avoid social desirability bias; we ensured that no interviewee was acquainted with the interviewer. Moreover, we phrased the questions as neutral and open-ended to encourage unbiased and detailed responses. For each interview, we created a non-judgemental atmosphere and clarified that there are only valid answers to our questions with valuable insights important to our research. We also explained to the interviewees that their responses would be anonymized and no links to their identity would be possible.
@@
Three of the authors coded each interview together. After familiarizing themselves with the transcripts (Step 1), the authors independently coded them (Step 2) to identify HPO methods applied by practitioners, to extract practitioners’ goals in HPO, and to reveal contextual factors that influence the interviewees’ decisions for HPO methods. We incorporated contextual factors to better understand influences on practitioners’ decisions for using HPO methods. The three authors independently read the transcripts, identified quotes relevant to this study, and labeled the quote with a name (i.e., a code) that expresses a potentially relevant HPO method, goal, or contextual factor. Each of the three authors coded the transcripts independently without a predefined coding scheme. After coding each interview independently, the authors presented their coding results to each other and discussed and harmonized their results in multiple iterations per interview and across interviews. In each iteration, the analysts aimed for mutually exclusive codes while ensuring comprehensive coverage. Through this explorative approach involving three analysts per interview, we aimed to reduce biases arising opinions of single analysts. The first coding iteration revealed 241 preliminary codes.
@@
After coding the transcripts of all 20 interviews, the analysis of the last 7 transcripts did not reveal additional HPO methods, goals, and contextual factors. We assume to have reached theoretical saturation (Fusch and Ness, 2015; Guest et al., 2006) after the first 13 interviews with 7 interviews confirming this assumption. These last seven interviews analyzed were from three practitioners from industry and four from academia who hold a master’s degree and are pursuing a PhD in the field of applied machine learning. Given the diverse backgrounds of the interviewees and the substantial number of confirmatory interviews, we considered the results sufficiently robust and moved on with an online survey to collect evidence for the external validity of the coding results.
@@ 3.2. Online Survey
The online questionnaire was structured into four sections: Introduction, HPO Methods and Goals, Contextual Factor Integration, and Demographics. In the Introduction section, we described the motivation for and the structure of the questionnaire. In Methods and Goals, we showed participants a matrix that listed all goals and HPO methods identified in the interview study. Participants were asked to select all pairs of HPO methods and goals to achieve specific goals. Because the main purpose of the survey was to collect evidence for the validity of the results from the interview study, study participants could only select HPO methods (i.e., manual tuning, grid search, random search, Bayesian optimization, and evolutionary algorithms) and goals (e.g., increase ML model understanding, decrease necessary computations, and decrease practitioner effort) extracted from the previous interview study. In a second question, we asked participants to indicate, for each selected pair of HPO method and goal, whether they feel to have successfully achieved each goal.
@@
Finally, in the Demographics section, we collected information about the participants to provide context to the data collected.
@@
To solicit participants for the online questionnaire, we contacted practitioners via email and promoted the study via social media platforms. We did not invite participants from the interview study.
In total, 166 participants filled out the first question of the HPO Methods and Goals section with 85 participants completing the whole section. Only 57 participants also completed the subsequent Contextual Factor Integration section. Of those, 8 participants did not provide demographic details, leading to a total of 49 participants completing the questionnaire. Most of these 49 participants worked in large organizations with more than 500 employees, including automotive companies, companies specializing in IT support and services, and universities. Table 2 shows more demographic details about the participants who completed the questionnaire.
@@
By analyzing the responses to the online questionnaire, we sought to learn how frequently practitioners tend to choose which HPO methods to pursue specific goals, given which contextual factors. To prepare the analysis, we discarded all responses from participants who did not complete the questionnaire. We only analyzed completed survey responses. We extracted the number of identical responses and related them to each other.
@@ 4. Results
The study participants (i.e., the interviewees and the survey participants) applied five principal HPO methods to pursue six goals, influenced by fourteen contextual factors (see sections 4.1.14.1.3). The study participants pursued their goals with different self-reported success (see section  4.2).
@@ 4.1.1. Hyperparameter Optimization Methods Used by Practitioners
The study participants applied five HPO methods: manual tuning, grid search, random search, Bayesian optimization, and evolutionary algorithms. Each HPO method was discussed by between two and eight interviewees. Most interviewees discussed one or two HPO methods in detail. [^2] We asked the interviewees to discuss a prominent recent example of how they used HPO methods, not to discuss all HPO methods they ever used. Most survey participants primarily used manual tuning, followed by grid search, Bayesian optimization, random search, and evolutionary algorithms (see Figure 1a). Academic participants tended to rely more on manual tuning and grid search compared to those from industry. Most survey participants stated to have used at least three HPO methods in their past ML projects; about 2% have even used at least five HPO methods. Only roughly 15% of the survey participants have used only a single HPO method (see Figure 1b).
@@
Even though literature indicates that Bayesian optimization yields better results in a shorter time than evolutionary algorithms, grid search, and random search (e.g., (Snoek et al., 2012; Turner et al., 2021)), practitioners tend to use seemingly inferior HPO methods. Practitioners seem to not only aim at finding hyperparameter configurations for optimal ML model performance but also pursue different goals.
@@ 4.1.2. Goals of Practitioners Pursued with Different HPO Methods
The goal comply with target audience refers to aligning practitioners’ choices with the expectations of their target audience. Three interviewees from academia stated they had decided on HPO methods to comply with the expectations of their target audiences regarding applied HPO methods and the resulting ML model. For example, two of the three interviewees described Bayesian optimization as uncommon in their research communities and felt obliged to explain it in their scientific publications. This would have required additional explanations of Bayesian optimization, even though the authors assumed the exact HPO method was not relevant to their scientific work. Therefore, they decided to use grid search as they assumed this HPO method to be well-known in their research communities.
@@
Extensive searches for optimal hyperparameter configurations in large search spaces typically require substantial computational resources. Necessary computations for HPO can be decreased by using an HPO method that requires fewer compute resources than other methods but is still sufficiently useful.
@@
In contrast to the goal described before (comply with target audience), decreasing necessary computations is not self-contained. Instead, it functions as a supplementary objective, generally addressed in conjunction with at least one goal and in response to contextual constraints. For example, practitioners may seek to decrease necessary computations to optimize the use of limited resources, enabling adequate improvements in ML model performance.
@@
Practitioners choose HPO methods to reduce overhead, for example, in terms of the additional time required to understand the HPO method or to integrate the HPO method into workflows. One industry interviewee reported that the sheer number of advanced HPO method made it challenging to select the most suitable method for their use case. Experiencing the paradox of choice (Schwartz, 2004), practitioners felt uncomfortable committing to a particular HPO method and tool and therefore opted for manual tuning instead.
To decrease their efforts in HPO, the interviewees applied grid search and manual tuning. In particular, practitioners stated to have applied manual tuning to avoid efforts related to setting up HPO tools in cluster infrastructures.
@@
Like decrease necessary computations, the goal decrease practitioner effort is not self-contained, and commonly used in combination with other goals. For example, study participants used HPO tools to enhance ML model understanding by iteratively refining manually defined sets of hyperparameter values. Those practitioners leveraged HPO tools to automate the reconfiguration of defined hyperparameter sets to accelerate manual tuning.
@@
Increasing ML model understanding refers to reaching the state where a practitioner can predict changes in an ML model’s behavior due to tuning hyperparameter values based on an understanding of the ML model’s inner workings. To increase their understanding of ML models, the interviewees reported having applied manual tuning. The interviewees claimed that manual tuning can improve their understanding of hyperparameter influences on ML models as they formulate hypotheses about hyperparameter influences on ML models and evaluate them immediately. The interviewees explained that they iteratively improve their ML model understanding by tuning hyperparameter values and testing their hypotheses. This dynamic between ML practitioner and ML model is researched as interactive ML. Interactive ML supports practitioners in actively exploring the model space and testing hypotheses about hyperparameter effects (Jiang et al., 2019). In addition, visual analytics tools help practitioners by providing graphical representations of model behavior and parameter interactions (e.g.,  (Spinner et al., 2019; Heyen et al., 2020)).
@@
The survey results indicate that all goals extracted from the interview study are also pursued by the survey participants (see Figure 2). More than 95% of the survey participants pursued the goal increase ML model performance and 75% aimed to achieve decrease necessary computations. 75% sought to decrease practitioner effort. 67% of the practitioners aimed to increase ML model understanding. The least pursued goals are satisfy requirements (63%) and comply with target audience (53%).
@@
Figure 3 shows how often the survey participants used HPO methods to reach the goals identified in the interview study. 67% of the survey participants tried to decrease necessary computations by using Bayesian optimization and 59% using evolutionary algorithms. About 50% of the participants tried to decrease the necessary computations by applying manual tuning. Random search and grid search were least often used, with 43% and 29%, respectively. Decreasing practitioner effort was of interest for less than 50% of the participants, with very similar responses for all HPO methods (yet a notable exception of manual tuning with only 31%). The survey participants primarily used manual tuning to enhance their comprehension of ML models. Interestingly, participants also tried to use Bayesian optimization twice as often as grid search and random search to achieve this goal despite its black-box nature. Participants from academia have only tried to satisfy requirements or be compliant with target audience half as often as participants from industry.
@@
Practitioners mostly aimed to satisfy requirements using evolutionary algorithms (45%), Bayesian optimization (35%), and manual tuning (37%). The study participants less often used grid search (24%) and random search (20%) to achieve this goal.
Comply with target audience was the least pursued goal for the participants. About 45% of the participants used evolutionary algorithms for this goal. We could not uncover differences between manual tuning and Bayesian optimization, with 30% of the participants using the respective method. Only random search and manual tuning were very seldom used to achieve this goal (about 15%).
In contrast to their academic counterparts, industry participants rarely used Bayesian optimization or evolutionary algorithms to increase model understanding. Participants from academia preferred to use grid search to reduce their effort. Additionally, industry practitioners more frequently relied on manual tuning to meet specific requirements.
It is evident that participants employed different HPO methods to pursue similar goals, making it difficult to establish a clear mapping between individual methods and goals. For example, while one interviewee selected manual tuning to increase ML model understanding, another favored Bayesian optimization to pursue the same goal.
@@
The ambiguous responses suggest that the choice of HPO methods is not simply explainable based on practitioners’ goals. For example, practitioners used different HPO methods to reach identical goals. This underscores the complexity of understanding why practitioners choose specific HPO methods. To better understand practitioners’ reasons for selecting HPO methods, contextual factors that influence practitioners in their decisions for HPO methods to reach goals seem important. By also illuminating contextual factors in combination with the use of HPO methods and goals to be achieved, the motives of practitioners to use HPO methods should be better understood.